From GlueXWiki
Jump to: navigation, search

December 5, 2013 FDC meeting


  1. Installation status [1], [2], [3], [4] (all)
  2. Engineering (Bill)
    • FDC survey [5]
    • Cathode scan [6]
  3. Electronics (Chris, Nick)
  4. Other


Participants: Bill, Dave, Nick, Chris, Simon, Beni, and Lubomir.


- The work on the mesh was finished just before Thanksgiving. LV cabling: over the weekend Vlad connected 11 out of 14 bundles for the cathodes (each 8 cables) at the detector side. The three remaining bundles are at the bottom of the FDC; in addition 24 cables for the anode LV and threshold have to be connected.

- On Monday there was an attempt to insert the FDC into the bore but it was difficult to move it. After some fixes done by Bill and some improvements, there was another attempt to insert it: this time the reflector holders at the top side were very close to BCAL, one of them broke at the epoxy after hitting a screw from a BCAL electronics box. The holder was re-glued and the top of the holders only at the top FDC side were filed taking out several mm. Finally on Tuesday the FDC together with the mesh were inserted in the magnet and surveyed there (see pictures above).

- Today the FDC was pulled/pushed out of the magnet and placed in a position for cabling. Before that openings in the platform were made to get access to the bottom side. Also scaffolds were install at the floor below the FDC and the mesh. Lubomir started working on the LV cables at the bottom side. The plan is to finish connecting the LV cables to the detector by Monday. Then we will connect them temporary to the LV panels and start testing. For that we will need the cooling system and Bill and Dave are going to connect it tomorrow.


- Bill: there are discrepancies in the survey results and Bill is going to work on the survey data; we are also waiting for the results of the survey done on Tuesday. Since the reflector holders only at one side hit the BCAL it means FDC is not on the same axis (by ~5 mm?), hopefully it is closer to the beam axis.

- We discussed how to scan the positions of the strips on the spare cathode. Lubomir obtained best resolution from the full package tests when decreasing the wire pitch by 10 microns which is unrealistic (1mm over the whole length!). We suspect it is the strip pitch that has to be modified and that's why this scan is very important. We want to measure the strip pitch along several lines (7 or 9) perpendicular to the strips. Bill will talk to the survey people about this.


- No meeting next week: Holiday party